Blaze Corp. applies overhead on the basis of direct labor hours. For the month of March, the company planned production of 10,000 units (80% of its production capacity of 12,500 units) and prepared the following budget:

Operating Levels
Overhead Budget 80%
Production in units 10,000
Standard direct labor hours 30,000
Budgeted overhead
Variable overhead costs
Indirect materials $ 30,000
Indirect labor 40,000
Power 8,000
Maintenance 3,000
Total variable costs 81,000
Fixed overhead costs
Rent of factory building 31,000
Depreciation—Machinery 45,000
Taxes and insurance 3,600
Supervisory salaries 16,400
Total fixed costs 96,000
Total overhead costs $ 177,000

During March, the company operated at 90% capacity (11,250 units), and it incurred the following actual overhead costs.

Overhead Costs
Indirect materials $ 30,000
Indirect labor 40,000
Power 9,000
Maintenance 3,810
Rent of factory building 31,000
Depreciation—Machinery 40,000
Taxes and insurance 4,350
Supervisory salaries 21,500
Total actual overhead costs $ 179,660


1. Compute the overhead controllable variance.
2. Compute the overhead volume variance.
3. Prepare an overhead variance report at the actual activity level of 9,000 units.

Required 1 Required 2 Required 3 Compute the overhead controllable variance. Classify as favorable or unfavorable. (Do not round intermediate calculations.) Controllable Variance Total actual overhead Flexible budget overhead Fixed Variable Total 0 Overhead controllable variance avorable
Required 1Required 2Required 3 Compute the overhead volume variance. Classify as favorable or unfavorable. (Do not round intermediate calculations.) Volume Variance otal budgeted fixed OH Total fixed overhead applied Volume variance avorable
Prepare an overhead variance report at the actual activity level of 9,000 units. Classify as favorable or unfavorable. (Do not round intermediate calculations.) BLAZE CORP Overhead Variance Report For Month Ended March 31 Expected production volume (80% of capacity 90% of capacity Production level achieved Volume variance Favorable Controllable Variance Flexible Budget Actual ResultsVariances Fav. / Unfav Variable overhead costs Indirect materials avorable Indirect labor avorable Power o variance Maintenance nfavorable otal variable costs avorable Fixed overhead costs Rent of factory building o variance Depreciation-Machinery avorable Taxes and insurance nfavorable Supervisory salaries nfavorable Unfavorable Total fixed costs Total overhead costs Favorable
Answer #1

Part 1

Controllable Variance

Total actual overhead

$179660

Flexible budget overhead

Variable

$91125

Fixed

96000

Total

187125

Overhead controllable variance

$7465

Favorable

Variable = 81000*90%/80%

Part 2

Volume Variance

Total budgeted fixed OH

$96000

Total fixed overhead applied

108000

Volume variance

$12000

Favorable

Standard rate = Budgeted fixed overhead/Budgeted direct labor hours = 96000/30000 = $3.20/hr

Fixed overhead applied = 33750 hours * $3.20/hr = $108000

(30000 hours / 80% *90% = 33750 hours)

Part 3

BLAZE CORP.

Overhead Variance Report

For Month Ended May 31

Expected production volume

80% of capacity

Production level achieved

90% of capacity

Volume variance

$12000

Favorable

Controllable Variance

Flexible Budget

Actual Results

Variances

Fav./Unfav.

Variable overhead costs:

Indirect materials

33750

30000

3750

Favorable

Indirect labor

45000

40000

5000

Favorable

Power

9000

9000

No variance

Maintenance

3375

3810

435

Unfavorable

Total variable costs

91125

82810

8315

Favorable

Fixed overhead costs:

Rent of factory building

40000

40000

No variance

Depreciation—Machinery

4350

4350

No variance

Supervisory salaries

16400

21500

5100

Unfavorable

Total fixed costs

60750

65850

5100

Unfavorable

Total overhead costs

151875

148660

$3215

Unfavorable


Indirect materials = 30000/80%*90% =33750

Indirect labor =40000/80%*90%

Power = 8000/80%*90% = 9000

Maintenance = 3000/80%*90% = 3375